A lawyer bills her clients $250 per hour of service. If a client's case requires 47 hours to complete, use proportion to calculate how much the client will owe the lawyer (excluding tax).

Answer by checkley71(8403)

1/250=47/x
1x=250*47
x=$11,750 is the total fee for 47 hours.